OpenStudy (anonymous):

y varies directly as to x. When x = 6 y = 2. Find y when x = 9.

OpenStudy (anonymous):

As y varies directly to x. So, y=kx we get, x×y=k If x=6 and y=2, x×y=k=6×2 so,k=12

OpenStudy (anonymous):

@mayankdevnani we answerd the same?

OpenStudy (anonymous):

No devnani

OpenStudy (anonymous):

x=6 and y=2, so 3y =x If x= 9, what will y be (y= 1/3x= 3 so in that case y=3